OpenStudy (anonymous):

2-(a-15/a+5) = (a^-5/a+5)

OpenStudy (anonymous):

Do you mean: \[2-\frac{a-15}{a+5}=\frac{a^{-5}}{a+5}\] If so, then: \[\frac{2(a+5)}{a+5}-\frac{a-15}{a+5}=\frac{a^{-5}}{a+5}\] \[\frac{a+25}{a+5}=\frac{a^{-5}}{a+5}\Rightarrow a+25=a^{-5}\] \[a^{-5}-a=25\] I'd need a calc to solve from here.